OpenStudy (anonymous):

Solve log2 16 – x = 7

OpenStudy (anonymous):

Is it 16 or is it (16-x)?

OpenStudy (anonymous):

16-x

OpenStudy (anonymous):

Solve log2 16 – x = 7 log2 (16 - x) = log2 (128) 16-x=128 x = -112 x=-3

OpenStudy (anonymous):

Okay...then it's \[2^{7}=16-x\]

OpenStudy (anonymous):

128 = 16 - x x = -112

OpenStudy (anonymous):

\[\log_{2}(16-x) =7\]?? Right?

OpenStudy (anonymous):

I am not understanding how the answer can b -3

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\log_{2}16-x=7 \]

OpenStudy (anonymous):

1st one or 2nd one?

OpenStudy (anonymous):

2nd

OpenStudy (anonymous):

If it's the 2nd (more likely) option then yes, it is -3.

OpenStudy (anonymous):

x = -112

OpenStudy (anonymous):

Corey how are you getting -3

OpenStudy (anonymous):

Here are the steps:

OpenStudy (anonymous):

I didn't solve for x at the end because it's simple algebra...but... x + 7 = 4 -7 -7 ____________ x = -3

OpenStudy (anonymous):

Hope the image helped somewhat.

OpenStudy (anonymous):

Fantastic! Thank you

OpenStudy (anonymous):

Not a problem!

OpenStudy (anonymous):

Don't forget to mark a Best Response. Take care.